OpenStudy (anonymous):

the average weekly earnings in dollars for various industries are listed below. Find the percentile rank of 653. 793, 653, 582, 703, 608, 734, 838, 534, 886, 925

OpenStudy (anonymous):

lets first start here , in order to get the correct percentile that you are required to get first you must arrange your data accodingly ,that is in most cases a continuois data 534,582,608,653,703,734,793,838,886 and 925 now that we have continually alligned the data it is now easy to tell that the company that we are looking for is number 4 4th(perc)=(n+1/sum 0f scores) =653+1/6423 =(0.1018216)*100 =10.18 there for 10.18 is the percentile rank of 653